# Site-Decorated Model for Unconventional Frustrated Magnets with Ultranarrow Phase Crossover and Spin Reversal Transition

### v1

@@ Abstract
The site-decorated Ising model is introduced to advance the understanding and physical realization of the recently discovered one-dimensional finite-temperature ultranarrow phase crossover in an external magnetic field, overcoming the complexity of the traditional bond-decorated models from geometric consideration. Furthermore, for higher-dimensional Ising models in the presence of an external magnetic field, while they remain unsolved, an exact solution about a novel spin-reversal transition—accessible by a slight change in temperature or the magnetic field, even in the weak field limit—is found to exist upon the site decoration. These results suggest a new route to energy-efficient applications in, e.g., data storage and processing, and call for materialization and device design with site decoration in, e.g., mixed $d$-$f$ compounds, optical lattices, or neural networks.
@@ I Introduction
Finding new states with desirable physical properties and energy-efficient ways of phase transition to these states is a central problem in many research fields ranging from condensed matter physics and materials science to quantum information and mircroelectronics. For example, magnetization reversal in a ferromagnet or ferrimagnet is a basic operation in data storage and processing, but the traditional magnetic-field-driven spin flipping could suffer from considerable coercive force in those materials [1]; the current-induced spin–orbit torque has emerged as a candidate for more effective approaches to magnetization manipulation [2]. On the other hand, new states with fast response abound in frustrated magnets as spin ice, spin glass, spin liquid, spin vortex and skyrmion, etc., [3, 4, 5]. To date, the conventional frustration, termed as geometrical frustration, describes a macroscopic ground-state degeneracy in the spin arrangement caused by competing spin-spin interactions  [6]. For example, to geometrically frustrate the interaction between two spins, couple them to one additional spin [Figs. 1(a)(c)]: Six spin configurations are degenerate when the three antiferromagnetic bonds of the resulting spin triangle are equally strong. This approach is referred to as bond decoration (i.e., decorating the bond between the first two spins). Bond decoration with strong geometric frustration has been widely used to explore advanced systems with novel phase changes [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], such as the recently discovered finite-temperature ultranarrow phase crossover (UNPC) in one-dimensional Ising model, where a finite-temperature phase transition is forbidden but now found approachable by UNPC [11, 12, 13].
The purpose of this paper is to introduce a site-decorated model for UNPC and phase transition, where the conventional geometric frustration is apparently absent [Figs. 1(b)(d)]. The idea was stimulated by the UNPC phenomenon driven by a hidden “half ice, half fire” (HIHF) state in a one-dimensional (1D) bond-decorated ferrimagnet [13, 14]. This exotic state has a macroscopic degeneracy, but as the phase boundary between excited states, it is always hidden in the ground-state phase diagram. The concept of atomic-scale mixed ice-fire states offers a new materials-design guide [15]. Although the ferrimagnet studied was still a bond-decorated system, the spin-spin interactions are not competing at all when $J$ is ferromagnetic [Figs. 1(a)(c)]. Such irrelevance of geometric frustration induced ground-state degeneracy suggests that it be likely and timely to bypass the complexity of bond decoration in search for unconventional frustrated magnets that are much easier to be synthesized and understood.
@@
Figure 1: Schematics of minimally decorated Ising models in one and two dimensions: (a)(c) bond decoration and (b)(d) site decoration. Green balls depict the original, type-a spins coupled by the interaction $J$ (red and orange bonds); gray balls depicts the decorated, type-b spins coupled to the backbone by the antiferromagnetic interaction $J_{b}<0$ (gray bonds). The spins’ magnetic moments are $\mu_{a}$ and $\mu_{b}$, respectively. The conventional geometric frustration is present [7, 8] and absent [13, 14] in the bond-decorated models (a)(c) for antiferromagnetic and ferromagnetic $J$, respectively. Such geometric frustration is clearly absent in the site-decorated models (b)(d).
@@
Here, using the Ising model, I show exactly that the site-decorated ferrimagnet displays a similar HIHF-driven UNPC at fixed finite temperature in one dimension. Furthermore, for the two- and three-dimensional (2D and 3D) site-decorated Ising models in an external magnetic field, while they remain unsolved, I am able to rigorously reveal a HIHF-driven spin reversal transition. The spin reversal is accessible by a slight change in temperature or the magnetic field, even in the weak field limit, suggesting a new route to energy-efficient applications in data storage and processing for example. Implications on material and device design with, e.g., mixed $d$-$f$ compounds [1, 15], are addressed. These results establish site decoration as a new, feasible avenue to the exploration of novel physical properties in unconventional frustrated magnets.
@@ II.1 One dimension
$H_{b}$ $=$ $-2J_{b}\sum_{i=1}^{N}\sigma_{i}b_{i}-h\mu_{b}\sum_{i=1}^{N}b_{i}.$ (2)
$H_{a}$ describes the backbone of the single chain with $\sigma_{i}=\pm 1$ (green balls, referred to as type-a spins) and $J>0$ the ferromagnetic interaction (red bonds). This is the model originally studied by Ising one century ago [16]. $H_{b}$ describes the decorated parts, where $b_{i}=\pm 1$ (gray balls, referred to as type-b spins) is coupled to the $i$th type-a spin with the antiferromagnetic interaction $02J_{b}<$ 0 (gray bonds). $h$ depicts the magnetic field, $\mu_{a}$ and $\mu_{b}$ the magnetic moments of type-a and type-b spins, respectively. The relationship of $\mu_{b}>\mu_{a}>0$ is used to represent ferrimagnetism—the model is invariant under the sign changes that satisfy $\mu_{a}\mu_{b}J_{b}<0$, e.g., $\mu_{a}=2$ and $\mu_{b}=-3$ in Sr3CuIrO6 [17, 18]. $N$ is the total number of the unit cell with $\sigma_{N+1}\equiv\sigma_{1}$, $b_{N+1}\equiv b_{1}$, viz., the periodic boundary condition.
In the calculation of the partition function $Z=\mathrm{Tr}\,e^{-\beta H}$ and the free energy per unit cell $f=-\frac{1}{N\beta}\ln Z$ in the thermodynamic limit $N\to\infty$, where $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the absolute temperature and $k_{\mathrm{B}}$ the Boltzmann constant, the type-b spins can be summed out exactly, resulting in $Z=e^{NA}\mathrm{Tr}\,e^{-\beta H_{\mathrm{eff}}}$ where
$A=\frac{1}{2}\left(\ln a+\ln b\right)$ (3)
with $a=2\cosh(\beta h\mu_{b}+2\beta J_{b})$ and $b=2\cosh(\beta h\mu_{b}-2\beta J_{b})$. $H_{\mathrm{eff}}$ is the effective Hamiltonian for the type-a spins:
@@
$h_{\mathrm{eff}}=h+\frac{1}{2\beta\mu_{a}}\left(\ln a-\ln b\right).$ (5)
$h_{\mathrm{eff}}$ is independent of $J$. Compared with the bond-decorated Ising model, $J$ is not normalized in the site-decorated model, which greatly simplifies the analysis.
Figure 2: Density plot of (a) $h_{\mathrm{eff}}$ and (b) $\langle\sigma_{i}\rangle$ in the $h$-$T$ plane for the 1D site-decorated Ising model. The black solid line in (a) is the $h_{\mathrm{eff}}=0$ contour line, i.e., the $T_{0}$ line defined in Equation (7), which is the $\langle\sigma_{i}\rangle=0$ contour line in (b) as the phase boundary between the red and purple regions with $\langle\sigma_{i}\rangle=\pm 1$, respectively. As shown in (b), at zero temperature, there are two phases separated at $h=h_{c}\equiv 2(-J_{b})/\mu_{a}$: (i) the ferrimagnetic state with all type-a spins down and all type-b spins up for $h<h_{c}$, and (ii) the ferromagnetic state with all spins up for $h>h_{c}$. The dashed line at $h=h_{f}\equiv 2(-J_{b})/\mu_{b}$, which is the phase boundary of two excited phases there (the ferromagnetic state and the opposite ferrimagnetic state with all type-a spins up and all type-b spins down), indicates where the half-ice, half-fire state (where all type-a spins are ordered up and all type-b spins are disordered) hides. Here $J_{b}=-1$, $\mu_{a}=1$, and $\mu_{b}=4/3$.
The order parameter that allows for fast and accurate identification of UNPC is the magnetization of the type-a spins given by $\langle\sigma_{i}\rangle=-\frac{\partial f}{h\,\partial\mu_{a}}=-\frac{\partial(f+A/\beta)}{\partial(h_{\mathrm{eff}}\mu_{a})}$ [13, 14]:
@@
which switches sign at the crossover temperature $T_{0}$ with the crossover width $2\delta T=2\left(\frac{\partial\langle\sigma_{i}\rangle}{\partial T}\right)^{-1}_{T=T_{0}}\propto e^{-2J/k_{\mathrm{B}}T_{0}}$. Apparently, $T_{0}$ can be determined by $h_{\mathrm{eff}}=0$ [10, 13, 14], yielding
$e^{4\beta J_{b}}=\frac{\sinh[\beta h(\mu_{b}-\mu_{a})]}{\sinh[\beta h(\mu_{b}+\mu_{a})]}.$ (7)
@@
It has a finite-temperature solution for $\mu_{b}>\mu_{a}$, $J_{b}<0$, and $0<h<2(-J_{b})/\mu_{a}$. As shown in Fig. 2(a) for $h_{\mathrm{eff}}$, Fig. 2(b) for $\langle\sigma_{i}\rangle$, and End Matter for $\langle b_{i}\rangle$, the essential physics of UNPC driven by the hidden half-ice (type-a spins are fully ordered), half-fire (type-b spins are fully disordered) state in the bond-decorated Ising model [14] has been reproduced in the site-decorated model, where $\langle\sigma_{i}\rangle$ flips from $+1$ to $-1$ with $T_{0}$ fixed and $2\delta T$ narrowed exponentially as $J$ increases. As $h$ increases, both $T_{0}$ and $\delta T$ decrease until $T_{0}=0$ and $02\delta T=$ 0 at $h=2(-J_{b})/\mu_{a}$. The largest $T_{0}$ as a function of $h$ takes place at the weak field limit $h\to 0$,
$k_{\mathrm{B}}T^{\mathrm{max}}_{0}=4J_{b}\left(\ln\frac{1-\mu_{a}/\mu_{b}}{1+\mu_{a}/\mu_{b}}\right)^{-1}.$ (8)
@@
Note that with the absolute zero field $h=0$, there is no sign change in $h_{\mathrm{eff}}$ for any finite temperature because $h_{\mathrm{eff}}\equiv 0$. This is consistent with the theorem that spontaneous UNPC does not exist in single-chain Ising models due to the spin-reversal symmetry associated with $h=0$ [11]. The minimal model for spontaneous UNPC is a decorated two-leg ladder Ising model [11, 12]. This abrupt change in the system’s behavior for $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
Figure 3: $T_{0}^{\mathrm{max}}$ (i.e., $T_{0}$ in the limit $h\to 0$) as a function of $\mu_{a}/\mu_{b}$. Here $J_{b}=-1$ and $\mu_{a}=1$. $T_{0}$ is independent of $J$.
@@
As shown in Fig. 4, $T^{\mathrm{max}}_{0}$ quickly increases as $\mu_{a}/\mu_{b}$ decreases, a feature desirable for low-field, high-temperature applications; however, the crossover width also increases as $2\delta T\propto e^{-2J/k_{\mathrm{B}}T_{0}}/h$ in the weak $h$ limit. The divergence of $2\delta T$ at $h=0$ for fixed $J$ agrees with the aforementioned theorem, which however can be overcome in higher dimensions, as proven below.
@@
Figure 4: The reversal of $\langle\sigma_{i}\rangle$ at $T_{0}$ as $h_{\mathrm{eff}}\to 0^{\pm}$ in the site-decorated square-lattice Ising model for $J=1$ with $T_{0}^{\mathrm{max}}<T_{c}^{*}\approx 2.269/k_{\mathrm{B}}$ (blue solid and dashed lines) and $J=0.88$ with $T_{0}^{\mathrm{max}}>T_{c}^{*}\approx 2/k_{\mathrm{B}}$ (red solid and dashed lines). Here $J_{b}=-1$, $\mu_{a}=1$, $\mu_{b}=4/3$ yielding $T_{0}^{\mathrm{max}}\approx 2.056/k_{\mathrm{B}}$.
II.2 Two dimension
@@
Although the Ising model in two dimension with nonzero magnetic field remains unresolved, it is appreciated that site decoration affects only the effective magnetic field in the effective Hamiltonian and $h_{\mathrm{eff}}=0$ at $T_{0}$. Equations (3)-(5), (7), and (8) can be directly applied to the 2D cases. Therefore, one can reach the following exact result as a function of $h$.
@@
For a site-decorated square lattice [Fig. 1(d)], at $T_{0}$ it reduces to a simple square lattice, which would undergo a paramagnetic-ferromagnetic phase transition at [19, 20]
@@
for $T_{0}<T_{c}^{*}$ and zero for $T_{0}>T_{c}^{*}$, where $\beta_{0}=1/(k_{\mathrm{B}}T_{0})$. Hence, the sign change of $h_{\mathrm{eff}}$ at $T_{0}$ results in a spin reversal transition. Recall that the sign change of $h_{\mathrm{eff}}$ can be achieved by a slight change in temperature or the magnetic field at the $T_{0}$ curve [the black line in Fig. 2(a)]. As shown in Fig. 4, there are two cases: (i) For $T_{0}^{\mathrm{max}}<T_{c}^{*}$ (blue solid and dashed lines), the spin reversal occurs in the whole region of $0<h<2(-J_{b})/\mu_{a}$. (ii) For $T_{0}^{\mathrm{max}}>T_{c}^{*}$ (red solid and dashed lines), the spin reversal has a threshold in $h$, whose corresponding $T_{0}$ equals $T_{c}^{*}$.
Note that $T_{c}^{*}$ is generally not the genuine Curie temperature $T_{c}$ of the system, which remains unresolved. $T_{c}^{*}$ is a value implied by $h_{\mathrm{eff}}=0$ and useful for determining which of the above two cases applies.
II.3 Three dimension
Likewise, since Equations (3)-(5), (7), and (8) can be directly applied to the site decoration in any dimension, the spin reversal transition as $h_{\mathrm{eff}}\to 0^{\pm}$ takes place in the site-decorated 3D Ising model in an external magnetic field as well with two cases similar to Fig. 4, although the exact expression of the spontaneous magnetization, like Equation (10) for a square lattice, is not available for the 3D Ising model.

### v2

# Site-Decorated Model for Unconventional Frustrated Magnets: Ultranarrow Phase Crossover and Spin Reversal Transition

@@ Abstract
The site-decorated Ising model is introduced to advance the understanding and experimental realization of the recently discovered one-dimensional finite-temperature ultranarrow phase crossover in an external magnetic field, while mitigating the geometric complexities of traditional bond-decorated models. Furthermore, although higher-dimensional Ising models in an external field remain unsolved, an exact solution for a novel spin-reversal transition—driven by an exotic, hidden ‘half-ice, half-fire’ state induced by site decoration—is derived. This transition, triggered by a slight variation in temperature or magnetic field even in the weak-field limit, offers a promising route toward energy-efficient applications such as data storage and processing. The results establish site decoration as a compelling new avenue for materials and device design, particularly in systems such as mixed $d$-$f$ compounds, optical lattices, and neural networks.
@@ I Introduction
Finding novel states with desirable physical properties and energy-efficient phase transitions is a central challenge in various research fields, including condensed matter physics, materials science, quantum information, and microelectronics [1]. For example, magnetization reversal in ferromagnets and ferrimagnets is crucial for data storage and processing, yet traditional field-driven spin flipping suffers from considerable coercive forces [2], while current-induced spin–orbit torque has emerged as a more effective alternative [3].
Frustrated magnets host a rich variety of unconventional states with rapid response dynamics, including spin ice, spin glass, spin liquid, spin vortex, and skyrmions [4, 5, 6]. Traditionally, frustration—termed geometric frustration—arises from competing spin-spin interactions, leading to macroscopic ground-state degeneracy [7]. A well-established approach to engineering such frustration is bond decoration, where an additional spin is introduced along a bond [Figs. 1(a)(c)] [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Notably, a finite-temperature ultranarrow phase crossover (UNPC) was recently discovered in one-dimensional (1D) bond-decorated Ising models with short-range interactions, both in the absence [13, 14] and presence [15] of an external magnetic field. While finite-temperature phase transitions have been theoretically forbidden in this fundamental 1D statistical mechanics system for a century [1, 16, 17], UNPC now enables them to be approached arbitrarily closely [13, 14, 15].
The purpose of this Letter is to introduce site decoration as an alternative paradigm for exotic phases, phase transitions, and UNPC, where additional spins are placed on lattice sites rather than bonds, thereby eliminating conventional geometric frustration [Figs. 1(b)(d)]. This idea was inspired by the UNPC phenomenon driven by a hidden half-ice, half-fire (HIHF) state [18] and a 1/3-ice, 2/3-fire state [15] in 1D bond-decorated ferrimagnets. These ice-fire states exhibit macroscopic degeneracy yet remain concealed in the ground-state phase diagram, emerging only as phase boundaries between excited states. Their existence suggests that atomic-scale mixed ice-fire states could serve as a guiding principle for materials design [19]. While these ferrimagnets were bond-decorated, their spin-spin interactions did not induce geometric frustration since $J$ was ferromagnetic [Figs. 1(a)(c)]. This key observation indicates that bypassing the complexities of bond decoration is both feasible and advantageous, paving the way for site-decorated frustrated magnets that are easier to synthesize and understand.
@@
Figure 1: Schematics of minimally decorated Ising models in one and two dimensions: (a)(c) bond decoration and (b)(d) site decoration. Green balls depict the backbone, type-a spins coupled by the interaction $J$ (red and orange bonds); gray balls depicts the decorated, type-b spins coupled to the backbone by the antiferromagnetic interaction $J_{ab}$ (gray bonds). The spins’ magnetic moments are $\mu_{a}$ and $\mu_{b}$, respectively. The conventional geometric frustration is present [9, 10] and absent [15, 18] in the bond-decorated models (a)(c) for antiferromagnetic and ferromagnetic $J$, respectively. Such geometric frustration is clearly absent in the site-decorated models (b)(d).
@@
Using the Ising model, I demonstrate exactly that the site-decorated ferrimagnet exhibits an HIHF-driven UNPC at a fixed finite temperature in one dimension. Furthermore, while the two- and three-dimensional (2D and 3D) site-decorated Ising models in an external magnetic field remain unsolved, I rigorously reveal a HIHF-driven spin reversal transition. This spin reversal can be induced by a slight change in temperature or magnetic field, even in the weak-field limit, suggesting a new pathway for energy-efficient applications in data storage and processing. Implications for material and device design, particularly in mixed $d$-$f$ compounds [2, 19], are also discussed. These findings establish site decoration as a novel and viable framework for exploring unique physical properties in unconventional frustrated magnets.
@@ II.1 One dimension
$H_{b}$ $=$ $-J_{ab}\sum_{i=1}^{N}\sigma_{i}b_{i}-h\mu_{b}\sum_{i=1}^{N}b_{i}.$ (2)
$H_{a}$ describes the backbone of the single chain with $\sigma_{i}=\pm 1$ (green balls, referred to as type-a spins) and $J>0$ the ferromagnetic interaction (red bonds). This is the model originally studied by Ising one century ago [16]. $H_{b}$ describes the decorated parts, where $b_{i}=\pm 1$ (gray balls, referred to as type-b spins) is coupled to the $i$th type-a spin with the antiferromagnetic interaction $J_{ab}<0$ (gray bonds). $h$ depicts the magnetic field, $\mu_{a}$ and $\mu_{b}$ the magnetic moments of type-a and type-b spins, respectively. The relationship of $\mu_{b}>\mu_{a}>0$ is used to represent ferrimagnetism—the model is invariant under the sign changes that satisfy $\mu_{a}\mu_{b}J_{ab}<0$, e.g., $\mu_{a}=2$ and $\mu_{b}=-3$ in Sr3CuIrO6 [20, 21]. $N$ is the total number of the unit cell with $\sigma_{N+1}\equiv\sigma_{1}$, $b_{N+1}\equiv b_{1}$, viz., the periodic boundary condition.
In the calculation of the partition function $Z=\mathrm{Tr}\,e^{-\beta H}$ and the free energy per unit cell $f=-\frac{1}{N\beta}\ln Z$ in the thermodynamic limit $N\to\infty$, where $\beta=1/(k_{\mathrm{B}}T)$ with $T$ being the absolute temperature and $k_{\mathrm{B}}$ the Boltzmann constant, the type-b spins can be summed out exactly for any dimension, resulting in $Z=e^{NA}\mathrm{Tr}\,e^{-\beta H_{\mathrm{eff}}}$ where
$A=\frac{1}{2}\left(\ln\boxed{+}+\ln\boxed{-}\right)$ (3)
with the shorthand notation ±$\boxed{\pm}=2\cosh(\beta h\mu_{b}\pm\beta J_{ab})$. $H_{\mathrm{eff}}$ is the effective Hamiltonian for the type-a spins:
@@
$h_{\mathrm{eff}}=h+\frac{1}{2\beta\mu_{a}}\left(\ln\boxed{+}-\ln\boxed{-}\right).$ (5)
$h_{\mathrm{eff}}$ is independent of $J$. Unlike in the bond-decorated Ising model, $J$ is not normalized in the site-decorated model, significantly simplifying the analysis.
Figure 2: Density plot of (a) $h_{\mathrm{eff}}$ and (b) $\langle\sigma_{i}\rangle$ in the $h$-$T$ plane for the 1D site-decorated Ising model. The black solid line in (a) is the $h_{\mathrm{eff}}=0$ contour line, i.e., the $T_{0}$ line defined in Equation (7), which is the $\langle\sigma_{i}\rangle=0$ contour line in (b) as the phase boundary between the red and purple regions with $\langle\sigma_{i}\rangle=\pm 1$, respectively. As shown in (b), at zero temperature, there are two phases separated at $h=h_{c}\equiv-J_{ab}/\mu_{a}$: (i) the ferrimagnetic state with all type-a spins down and all type-b spins up for $h<h_{c}$, and (ii) the ferromagnetic state with all spins up for $h>h_{c}$. The dashed line at $h=h_{f}\equiv-J_{ab}/\mu_{b}$, which is the phase boundary of two excited phases there (the ferromagnetic state and the opposite ferrimagnetic state with all type-a spins up and all type-b spins down), indicates where the half-ice, half-fire state (where all type-a spins are ordered up and all type-b spins are disordered) hides. Here $J_{ab}=-2$, $\mu_{a}=1$, and $\mu_{b}=4/3$.
Figure 3: $T_{0}^{\mathrm{max}}$ (i.e., $T_{0}$ in the limit $h\to 0$) as a function of $\mu_{a}/\mu_{b}$. Here $J_{ab}=-2$ and $\mu_{a}=1$. $T_{0}$ is independent of $J$.
The order parameter that allows for fast and accurate identification of UNPC is the magnetization of the type-a spins given by $\langle\sigma_{i}\rangle=-\frac{\partial f}{h\,\partial\mu_{a}}$ [15, 18]:
@@
which switches sign at the crossover temperature $T_{0}$. Then, according to Equation (6), $T_{0}$ can be determined by setting $h_{\mathrm{eff}}=0$, yielding the critical equation:
$\tanh(-\beta J_{ab})=\frac{\tanh(\beta h\mu_{a})}{\tanh(\beta h\mu_{b})},$ (7)
$T_{0}$ elegantly links four parameters of the Hamiltonian, while the remaining parameter, $J$, exponentially narrows the crossover width $2\delta T=2\left(\frac{\partial\langle\sigma_{i}\rangle}{\partial T}\right)^{-1}_{T=T_{0}}\propto e^{-2J/k_{\mathrm{B}}T_{0}}$. Equation (7) has a finite-temperature solution for $|\mu_{a}|<|\mu_{b}|$, $\mu_{a}\mu_{b}J_{ab}<0$, and $0<h<h_{c}$, where $h_{c}\equiv|J_{ab}/\mu_{a}|$.
@@
As shown in Fig. 2(a) for $h_{\mathrm{eff}}$, Fig. 2(b) for $\langle\sigma_{i}\rangle$, and End Matter for $\langle b_{i}\rangle$, the essential physics of UNPC—driven by the hidden half-ice (type-a spins are fully ordered), half-fire (type-b spins are fully disordered) state in the bond-decorated Ising model [18]—has been reproduced in the site-decorated model. In this framework, $\langle\sigma_{i}\rangle$ flips from $+1$ to $-1$ at a fixed $T_{0}$, while $2\delta T$ narrows exponentially as $J$ increases.
As $h$ increases, both $T_{0}$ and $\delta T$ decrease, reaching $T_{0}=0$ and $02\delta T=$ 0 at $h_{c}$. The maximum $T_{0}$ as a function of $h$ occurs at the weak field limit $h\to 0$,
$k_{\mathrm{B}}T^{\mathrm{max}}_{0}=2J_{ab}\left(\ln\frac{1-\mu_{a}/\mu_{b}}{1+\mu_{a}/\mu_{b}}\right)^{-1}.$ (8)
@@
Note that with the absolute zero field ($h=0$), there is no sign change in $h_{\mathrm{eff}}$ because $h_{\mathrm{eff}}=0$ at any finite temperature. This is consistent with the theorem that spontaneous UNPC does not exist in single-chain Ising models due to the spin-reversal symmetry associated with $h=0$ [13]. The minimal model for spontaneous UNPC is a decorated two-leg ladder Ising model [13, 14]. This abrupt change in the system’s behavior between $h=0$ and $h\to 0$ is rare and can be used to measure the absolute zero magnetic field.
@@
As shown in Fig. 4, $T^{\mathrm{max}}_{0}$ rapidly increases as $\mu_{a}/\mu_{b}$ decreases, a feature desirable for low-field, high-temperature applications. However, the crossover width also increases as $2\delta T\propto e^{-2J/k_{\mathrm{B}}T_{0}}/h$ in the weak $h$ limit. The divergence of $2\delta T$ at $h=0$ for fixed $J$ agrees with the aforementioned theorem; however, this limitation can be overcome in higher dimensions, as proven below.
@@
Figure 4: The reversal of $\langle\sigma_{i}\rangle$ at $T_{0}$ as $h_{\mathrm{eff}}\to 0^{\pm}$ in the site-decorated square-lattice Ising model for $J=1$ with $T_{0}^{\mathrm{max}}<T_{c}^{*}\approx 2.269/k_{\mathrm{B}}$ (blue solid and dashed lines) and $J=0.88$ with $T_{0}^{\mathrm{max}}>T_{c}^{*}\approx 2/k_{\mathrm{B}}$ (red solid and dashed lines). Here $J_{ab}=-2$, $\mu_{a}=1$, $\mu_{b}=4/3$ yielding $T_{0}^{\mathrm{max}}\approx 2.056/k_{\mathrm{B}}$.
II.2 Higher dimensions
@@
Although 2D and 3D Ising models with a nonzero magnetic field remains unresolved, it has been proven for any dimension that site decoration yields $h_{\mathrm{eff}}=0$ at $T_{0}$ as a function of $h$ [the black line in Fig. 2(a)], while keeping $J$ unchanged. Equations (3)-(5), (7), and (8) are directly applicable to the 2D and 3D cases. Consequently, the following exact result can be derived as a function of $h$.
@@
For a site-decorated square lattice [Fig. 1(d)], the system reduces to a simple square lattice with $h_{\mathrm{eff}}=0$ at $T_{0}$, which would undergo a spontaneous paramagnetic-ferromagnetic phase transition at [22, 23]
@@
for $T_{0}<T_{c}^{*}$ and zero for $T_{0}>T_{c}^{*}$, where $\beta_{0}=1/(k_{\mathrm{B}}T_{0})$.
Hence, the sign change of $h_{\mathrm{eff}}$ at $T_{0}$ induces a spin reversal transition. Recall that this sign change can be triggered by a slight variation in temperature or magnetic field along the $T_{0}$ curve [the black line in Fig. 2(a)]. As shown in Fig. 4, two cases arise: (i) If $T_{0}^{\mathrm{max}}<T_{c}^{*}$ (blue solid and dashed lines), spin reversal occurs throughout the entire region $0<h<h_{c}$. (ii) If $T_{0}^{\mathrm{max}}>T_{c}^{*}$ (red solid and dashed lines), spin reversal has a threshold in $h$, where the corresponding $T_{0}$ equals $T_{c}^{*}$.
Note that $T_{c}^{*}$ is generally not the true Curie temperature $T_{c}$ of the system, which remains unresolved. Instead, $T_{c}^{*}$ is a value derived for $h_{\mathrm{eff}}=0$ and serves as the criterion for determining which of the two cases applies.
Similarly, the spin reversal transition at $h_{\mathrm{eff}}=0$ ($h\neq 0$) also occurs in other site-decorated 2D and 3D Ising models under an external magnetic field. This transition follows two cases analogous to Fig. 4, with exact results calculable for the anisotropic rectangular, anisotropic triangular [25], honeycomb, and Kagomé [26] lattices, where the exact expression for spontaneous magnetization—akin to Equation (10) for a square lattice—is known.